# LARES: Latent Reasoning for Sequential Recommendation

### v1

@@ 3.4.4. Influence of Reasoning Steps on Inference Time
To systematically assess the impact of reasoning steps on inference efficiency, we conduct a comparative analysis between LARES with varying reasoning steps and SASRec under equivalent computational budgets (FLOPs) on the Instrument dataset. LARES uses 2-layer pre-blocks and core-blocks, and the batch size is set to 1024 for both. As presented in Table 4 (where item embeddings are excluded from parameter calculations), LARES achieves higher computational complexity while maintaining identical parameter sizes through increased reasoning steps. For instance, with 4 reasoning steps, LARES attains an effective depth of 2+2×4 layers, comparable to a 10-layer SASRec architecture. The additional inference latency introduced by LARES remains moderate—approximately 10% higher than SASRec at comparable FLOPs. This computational overhead is acceptable in consideration of LARES’s substantial performance gains, which improve SASRec’s baseline by an average of 20% on Instrument as shown in Figure 2. Our findings demonstrate that LARES offers an advantageous trade-off between computational cost and model performance, suggesting strong potential for practical deployment.
@@ 4.2. Reasoning Models
Recent advances in test-time scaling (DeepSeek-AI et al., 2025; Chen et al., 2025a, b) have shifted the research focus in GenAI from large language models (LLMs) to large reasoning models (LRMs). LRMs excel in complex reasoning tasks through deep thinking capabilities, as demonstrated by the powerful reasoning systems like OpenAI-o1 and DeepSeek-R1. These models employ explicit long Chain-of-Thought mechanisms to generate extensive reasoning tokens before producing final answers (Yang et al., 2024; Besta et al., 2024; Yang et al., 2025). However, their reliance on explicit reasoning poses some critical challenges, including excessive memory demands from long context windows and limited expressive power due to discrete language space constraints. To address these limitations, recent studies (Hao et al., 2024; Chen et al., 2024; Geiping et al., 2025; Xu et al., 2025a) have introduced latent reasoning models that perform implicit reasoning in continuous latent spaces, achieving greater efficiency. In recommender systems, some efforts have adapted reasoning mechanisms for sequential recommendation models (Tang et al., 2025; Zhang et al., 2025). For instance, ReaRec autoregressively generates latent reasoning tokens to refine user representations. STREAM-Rec integrates slow-thinking paradigms with TIGER-style generative recommenders by producing annotated reasoning tokens before final recommendation semantic tokens. In contrast, we propose LARES, a novel recurrent-depth latent reasoning framework for sequential recommendation. Unlike prior work, LARES iteratively refines all input tokens at each reasoning step. Additionally, it is seamlessly compatible with existing sequential recommendation models, further enhancing their performance.

### v2

@@ 3.4.4. Influence of Reasoning Steps on Inference Time
To systematically assess the impact of reasoning steps on inference efficiency, we conduct a comparative analysis between LARES with varying reasoning steps and SASRec under equivalent computational budgets (FLOPs) on the Instrument dataset. LARES uses a 2-layer pre-block and core-block, and the batch size is set to 1024 for both. As presented in Table 4 (where item embeddings are excluded from parameter calculations), LARES achieves higher computational complexity while maintaining identical parameter sizes through increased reasoning steps. For instance, with 4 reasoning steps, LARES attains an effective depth of 2+2×4 layers, comparable to a 10-layer SASRec architecture. The additional inference latency introduced by LARES remains moderate—approximately 10% higher than SASRec at comparable FLOPs. This computational overhead is acceptable in consideration of LARES’s substantial performance gains, which improve SASRec’s baseline by an average of 20% on Instrument as shown in Figure 2. Our findings demonstrate that LARES offers an advantageous trade-off between computational cost and model performance, suggesting strong potential for practical deployment.
@@ 4.2. Reasoning Models
Recent advances in test-time scaling (DeepSeek-AI et al., 2025; Chen et al., 2025a, b) have shifted the research focus in GenAI from large language models (LLMs) to large reasoning models (LRMs). LRMs excel in complex reasoning tasks through deep thinking capabilities, as demonstrated by the powerful reasoning systems like OpenAI-o1 and DeepSeek-R1. These models employ explicit long Chain-of-Thought mechanisms to generate extensive reasoning tokens before producing final answers (Yang et al., 2024; Besta et al., 2024; Yang et al., 2025). However, their reliance on explicit reasoning poses some critical challenges, including excessive memory demands from long context windows and limited expressive power due to discrete language space constraints. To address these limitations, recent studies (Hao et al., 2024; Chen et al., 2024; Geiping et al., 2025; Xu et al., 2025a) have introduced latent reasoning models that perform implicit reasoning in continuous latent spaces, achieving greater efficiency. In recommender systems, some efforts have adapted reasoning mechanisms for sequential recommendation models (Tang et al., 2025; Zhang et al., 2025). For instance, ReaRec (Tang et al., 2025) autoregressively generates latent reasoning tokens to refine user representations. STREAM-Rec (Zhang et al., 2025) integrates slow-thinking paradigms with TIGER-style generative recommenders by producing annotated reasoning tokens before final recommendation semantic tokens. In contrast, we propose LARES, a novel recurrent-depth latent reasoning framework for sequential recommendation. Unlike prior work, LARES iteratively refines all input tokens at each reasoning step. Additionally, it is seamlessly compatible with existing sequential recommendation models, further enhancing their performance.